A rental car agency charges $16 per day plus 13 cents per mile to rent a certain car. Another agency charges $21 per day plus 7 cents per mile to rent the same car. How many miles will have to be driven for the cost of a car from the first agency to equal the cost of a car from the second agency?(round answer to the nearest hundreth of a mile.)Express the problems as a system of linear equations and solve using the method of your choice. Please help me. I have no clue!
Answer by checkley75(3666)

1600+13m=2100+7m
13m-7m=2100-1600
6m=500
m=500/6
m=83.33 miles.
This says the the two fees are the same asfter 83.33 miles.
proof
1600+13*83.33=2100+7*83.33
1600+1083.33=2100+583.33
2683.33=2683.33